# A Tutorial on Beyond-Diagonal Reconfigurable Intelligent Surfaces: Modeling, Architectures, System Design and Optimization, and Applications

### v1

@@ V-C Wideband Effect
Frequency dependence is an intrinsic phenomenon in circuits, where the response of inductance and capacitance varies with the frequency of an applied signal. Therefore, the response of RIS, i.e., its impedance/admittance/scattering matrices of the reconfigurable impedance network, is naturally frequency dependent [93]. In narrowband communications, since the bandwidth of signals is much less than the central frequency, it is reasonable to approximately assume the response of RIS to be frequency independent. However, when it comes to wideband communications, the frequency dependence in RIS is not negligible. To show how significant is the frequency dependence in RIS-aided wideband communication systems, there have been a few representative works on the wideband modeling, optimization, and implementation of D-RIS [19, 94, 95]. Nevertheless, these works model the wideband effect of D-RIS by characterizing the frequency dependence of the individual entries of its scattering matrix. This does not work for BD-RIS with interconnected elements, and thus coupled entries in its scattering matrix, making it difficult to accurately yet efficiently capture the frequency dependence in BD-RIS. In addition, there have been a few works on applying BD-RIS in wideband systems [96, 97], while they primarily assume a frequency-independent BD-RIS model, which is essentially not consistent with the behavior of BD-RIS in wideband scenarios. To establish a physically consistent BD-RIS model suitable for wideband systems, [98] and [99] have proposed to model the frequency dependence of the individual reconfigurable admittance component based on specific circuit designs. Below, we will briefly revisit how to model the frequency dependence in BD-RIS.
Figure 21: An example of a 6-element BD-RIS with group-connected architectures and the equivalent lossless circuit for each admittance component.
@@
Figure 22: The susceptance $\Im\{Y_{m,m^{\prime}}\}$ as a function of (a) frequency and (b) the value of susceptance at the central frequency $f_{\mathsf{c}}=2.4$ GHz with $L_{1}=2.5$ nH, $L_{2}=0.7$ nH, and $C\in[0.2,3]$ pF for a practical varactor diode [98].
(a) Group-connected BD-RIS with and without wideband modeling
(b) Forest-connected BD-RIS with and without wideband modeling
Figure 23: Average rate versus transmit power $P$ with BD-RIS having different reciprocal architectures ($M=48$, $\bar{M}\in\{1,3,6\}$). The legend “WM” is short for wideband modeling; “GC” is short for group-connected; “FC” is short for forest-connected. For both “GC” and “FC” architectures, the case of $\bar{M}=1$ refers to the D-RIS [98].
In reciprocal BD-RIS architectures, each reconfigurable admittance component is modeled as a lumped circuit consisting of inductors $L_{1}$, $L_{2}$, and one tunable capacitor $C$, as illustrated in Fig. 21333Here we consider a lossless model with $R=0$ in Fig. 18 to focus purely on the modeling and analysis of frequency dependence in BD-RIS.. Accordingly, the admittance of this circuit is a function of $C$ and the angular frequency $\omega=2\pi f$ for the incident signals, that is
@@
where $Y_{m,m}$ boils down to the admittance component $Y_{m}$ connected to ground. We observe from (100) that, the values of $Y_{m,m^{\prime}}$ at different frequencies are linked to each other by a common capacitor $C$. The relationship can be numerically illustrated in Fig. 22, based on a wideband system with central frequency $f_{\mathsf{c}}=2.4$ GHz and a practical varactor diode SMV1231-079 [98]. Results in Fig. 22(a) show that within some practical bandwidth for wideband signals, the susceptance $\Im\{Y_{m,m^{\prime}}\}$ can be regarded as a linear function of frequency. This further implies that $\Im\{Y_{m,m^{\prime}}\}$ varies approximately linearly with the susceptance at central frequency, as illustrated in Fig. 22(b). This motivates possible simplifications to (100) to benefit BD-RIS optimization. The impact of having frequency-dependent BD-RIS matrices on the performance of a SISO orthogonal frequency division multiplexing (OFDM) system is shown in Fig. 23. Results show that, with increasing circuit complexity (more admittance components) in BD-RIS architectures, there is an increasing performance gap between taking into account the wideband modeling when designing BD-RIS or not. This is attributed to the more significant variation of BD-RIS matrices between subcarriers, such that ignoring this variation will cause non-negligible performance loss.
@@
[97] Group-Connected
@@
The above modeling is again applicable to all reciprocal BD-RIS architectures as illustrated in Section III-C. Intuitively, the higher the circuit complexity (the larger the required number of admittance components) in BD-RIS architectures, the more significant the impact of wideband modeling at BD-RIS. In addition, [100] has proposed to model and analyze the frequency dependence of non-reciprocal BD-RIS with constraint $\mathcal{T}_{\mathsf{non-diag}}$, by individually modeling the frequency-dependent scattering parameter of each reconfigurable element. Nevertheless, due to the lack of a unified circuit topology illustration as in reciprocal architectures, the wideband modeling and optimization for more general non-reciprocal BD-RIS architectures still remain unexplored. For clarity, the BD-RIS literature on wideband modeling and optimization has been summarized in Table VII.
V-D Mutual Coupling Effect
@@
Mutual coupling, which refers to the electromagnetic interaction between antenna elements in an array, is an important factor in array design. Mutual coupling is in general inversely proportional to the spacing between elements: the smaller the inter-element spacing, the stronger the mutual coupling. In most existing RIS literature, the mutual coupling is ignored by assuming sufficiently large inter-element spacing. However, this strong assumption cannot be easily achieved in practice, given that RIS usually consists of numerous densely spaced elements within a limited aperture to provide sufficient beamforming gain [8, 101]. There have been existing works on the modeling and analysis of the mutual coupling between D-RIS elements [18, 102, 103], which have been recently extended to BD-RIS-aided channels [104, 105, 106]. Below, we will briefly revisit the modeling of mutual coupling aware BD-RIS-aided channels, and how the existence of mutual coupling impacts the system performance.
@@
Consider a BD-RIS-aided MIMO system consisting of an $N$-antenna transmitter, an $M$-element BD-RIS, and an $N_{\mathsf{r}}$ receiver, as also illustrated in Section IV-B. To capture explicitly the mutual coupling at BD-RIS, [35, 104, 18, 103] have made physics-consistent assumptions and derived the following three equivalent wireless channel models [^4] Note that the scattering matrices $\mathbf{S}_{RT}$, $\mathbf{S}_{RI}$, and $\mathbf{S}_{IT}$ in (101a) are not one-to-one mapped to the channels $\mathbf{H}_{RT}$, $\mathbf{H}_{RI}$, and $\mathbf{H}_{IT}$ in the widely adopted model $\mathbf{H}(\mathbf{\Theta})=\mathbf{H}_{RT}+\mathbf{H}_{RI}\mathbf{\Theta}\mathbf{H}_{IT}$. The detailed analysis can be found in [35, 107, 108].
@@
where $\mathbf{S}_{RT}\in\mathbb{C}^{N_{\mathsf{r}}\times N}$, $\mathbf{Z}_{RT}\in\mathbb{C}^{N_{\mathsf{r}}\times N}$, and $\mathbf{Y}_{RT}\in\mathbb{C}^{N_{\mathsf{r}}\times N}$, respectively, denote the transmission scattering, impedance, and admittance matrices from the transmitter to receiver; $\mathbf{S}_{RI}\in\mathbb{C}^{N_{\mathsf{r}}\times M}$, $\mathbf{Z}_{RI}\in\mathbb{C}^{N_{\mathsf{r}}\times M}$, and $\mathbf{Y}_{RI}\in\mathbb{C}^{N_{\mathsf{r}}\times M}$, respectively, denote the transmission scattering, impedance, and admittance matrices from RIS to the receiver; $\mathbf{S}_{IT}\in\mathbb{C}^{M\times N}$, $\mathbf{Z}_{IT}\in\mathbb{C}^{M\times N}$, and $\mathbf{Y}_{IT}\in\mathbb{C}^{M\times N}$, respectively, denote the transmission scattering, impedance, and admittance matrices from the transmitter to RIS; $\mathbf{S}_{II}\in\mathbb{C}^{M\times M}$, $\mathbf{Z}_{II}\in\mathbb{C}^{M\times M}$, and $\mathbf{Y}_{II}\in\mathbb{C}^{M\times M}$, respectively, denote the mutual coupling scattering, impedance, and admittance matrices at RIS. Based on the derivations in [35] and [104], these terms are related to one another as follows
$\mathbf{S}_{RT}$ $=\frac{\mathbf{Z}_{RT}}{2Z_{0}}-\frac{\mathbf{Z}_{RI}}{2Z_{0}}(\mathbf{Z}_{II}+Z_{0}\mathbf{I}_{M})^{-1}\mathbf{Z}_{IT},$ (102a)
$\mathbf{S}_{RI}$ $=\mathbf{Z}_{RI}(\mathbf{Z}_{II}+Z_{0}\mathbf{I}_{M})^{-1},$ (102b)
$\mathbf{S}_{IT}$ $=(\mathbf{Z}_{II}+Z_{0}\mathbf{I}_{M})^{-1}\mathbf{Z}_{IT},$ (102c)
$\mathbf{S}_{II}$ $=(\mathbf{Z}_{II}+Z_{0}\mathbf{I}_{M})^{-1}(\mathbf{Z}_{II}-Z_{0}\mathbf{I}_{M}),$ (102d)
$\mathbf{Y}_{RT}$ $=\frac{1}{Z_{0}^{2}}(-\mathbf{Z}_{RT}+\mathbf{Z}_{RI}\mathbf{Z}_{II}^{-1}\mathbf{Z}_{IT}),$ (102e)
$\mathbf{Y}_{RI}$ $=-\frac{\mathbf{Z}_{RI}\mathbf{Z}_{II}^{-1}}{Z_{0}},\mathbf{Y}_{IT}=-\frac{\mathbf{Z}_{II}^{-1}\mathbf{Z}_{IT}}{Z_{0}},\mathbf{Y}_{II}=\mathbf{Z}_{II}^{-1}.$ (102f)
@@
where $\mathbf{\Theta}_{\mathsf{eff}}=(\mathbf{I}_{M}-\mathbf{\Theta}\mathbf{S}_{II})^{-1}\mathbf{S}_{II}$, one can also obtain a beyond-diagonal matrix $\mathbf{\Theta}_{\mathsf{eff}}$ for D-RIS. However, the BD-RIS terminology is used to describe the scattering matrix $\mathbf{\Theta}$. In this sense, within the BD-RIS, a kind of tunable coupling is artificially introduced between ports of the reconfigurable impedance network, thereby generating $\mathbf{\Theta}$ itself that is not limited to be diagonal.
Remark 9: Another interpretation of BD-RIS has been recently presented in [109], whose main idea is that an $M$-element BD-RIS-aided channel can be represented as a channel involving an effective D-RIS matrix $\tilde{\mathbf{\Theta}}\in\mathbb{C}^{\tilde{M}\times\tilde{M}}$, where $\tilde{M}$ denotes the number of reconfigurable admittance components in BD-RIS. Specifically, the channel model can be described as
@@
The impedance matrix $\mathbf{Z}_{II}$ characterizes the mismatching and mutual coupling at the RIS elements. Specifically, the diagonal entries of $\mathbf{Z}_{II}$ refer to the self impedance and the off-diagonal entries refer to the mutual coupling depending on the inter-element spacing. Most existing literature assumes perfect matching and no mutual coupling at RIS elements, which is mathematically described as $\mathbf{Z}_{II}=Z_{0}\mathbf{I}_{M}$. This assumption makes $\mathbf{S}_{II}=\mathbf{0}_{M\times M}$ in (102) and the wireless channel $\mathbf{H}_{\mathsf{S}}(\mathbf{\Theta})$ in (101) a linear function of $\mathbf{\Theta}$, as widely assumed in most existing RIS literature. However, as discussed above, it is difficult to achieve such a strong assumption in practice. That is, the off-diagonal entries of $\mathbf{Z}_{II}$ are generally nonzero, and can be modeled as functions of inter-element spacing [103, 18]. Specifically, [103] has modeled each off-diagonal entry of $\mathbf{Z}_{II}$ assuming isotropic radiators, that is
${}_{m,m^{\prime}}$ $=-\frac{[\mathbf{Z}_{II}]_{m,m}}{\jmath\kappa_{0}d|m-m^{\prime}|}e^{-\jmath\kappa_{0}d|m-m^{\prime}|},$ (105)
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\forall m\neq m^{\prime},m,m^{\prime}\in\mathcal{M},$
where $\kappa_{0}=\frac{2\pi}{\lambda}$ denotes the wave number, $d$ is the inter-element spacing, and $[\mathbf{Z}_{II}]_{m,m}$, $\forall m\in\mathcal{M}$ are set to identical constants. Alternatively, [18] has modeled each off-diagonal entry of $\mathbf{Z}_{II}$ assuming thin wire dipoles with radius $r$ and length $\iota$. Then, the mutual impedance between dipoles $m$ and $m^{\prime}$ with center-point coordinates $(m_{x},m_{y},m_{z})$ and $(m^{\prime}_{x},m^{\prime}_{y},m^{\prime}_{z})$ is computed as
${}_{m,m^{\prime}}$ $=\int_{m_{z}-\frac{\iota}{2}}^{m_{z}+\frac{\iota}{2}}\int_{m^{\prime}_{z}-\frac{\iota}{2}}^{m^{\prime}_{z}+\frac{\iota}{2}}\frac{\jmath\eta_{0}}{4\pi\kappa_{0}}\bigg{(}\frac{(n_{z}-n_{z}^{\prime})^{2}}{d_{m,m^{\prime}}^{2}}$ (106)
×$~{}\times\big{(}\frac{3}{d_{m,m^{\prime}}^{2}}+\frac{\jmath 3\kappa_{0}}{d_{m,m^{\prime}}}-\kappa_{0}^{2}\big{)}-\frac{\jmath\kappa_{0}+d_{m,m^{\prime}}^{-1}}{d_{m,m^{\prime}}}+\kappa_{0}^{2}\bigg{)}$
×$~{}\times\frac{e^{-\jmath\kappa_{0}d_{m,m^{\prime}}}}{d_{m,m^{\prime}}}\frac{\sin(\kappa_{0}(\frac{\iota}{2}-|n_{z}^{\prime}-m^{\prime}_{z}|))}{\sin^{2}(\kappa_{0}\frac{\iota}{2})}$
×$~{}\times\sin(\kappa_{0}(\frac{\iota}{2}-|n_{z}-m_{z}|))dn_{z}^{\prime}dn_{z},\forall m,m^{\prime}\in\mathcal{M},$
where $\eta_{0}=377$ $\Omega$ denotes the free-space impedance, and $d_{m,m^{\prime}}$ denotes the distance between the dipoles $m$ and $m^{\prime}$, i.e., $d_{m,m^{\prime}}=\sqrt{(d_{m,m^{\prime}}^{x,y})^{2}+(n_{z}^{\prime}-n_{z})^{2}}$, $d_{m,m^{\prime}}^{x,y}=\sqrt{(m_{x}-m^{\prime}_{x})^{2}+(m_{y}-m^{\prime}_{y})^{2}}$ for $m\neq m^{\prime}$ and $d_{m,m}^{x,y}=r$.
@@
Figure 24: Channel gain versus the number of BD-RIS elements ($M=16$, $\bar{M}=4$ for group-connected architecture). The markers “FC”, “GC”, “SC” respectively refer to fully-connected, group-connected, and single-connected (D-RIS); “w/ MC” refers to the case accounting for mutual coupling at BD-RIS and “w/o MC” refers to the case where the off-diagonal entries of $\mathbf{Z}_{II}$ are forced to zero [104].
@@
To evaluate how the existence of mutual coupling impacts the system performance, [104] has primarily focused on a group/fully-connected BD-RIS-aided SISO system and proposed an iterative algorithm to design the BD-RIS impedance matrix. Simulation results in Fig. 24 demonstrate that, under purely LoS channels, the performance gap between BD-RIS and D-RIS increases with decreasing inter-element distance since the stronger mutual coupling is better exploited by BD-RIS architectures. This indicates that BD-RIS is suitable for compact deployment with small inter-element spacing. In addition, [105] has proposed decoupling networks as a solution to compensate for the mutual coupling at BD-RIS. The decoupling network can transform the system into a structure equivalent to a system without mutual coupling. As such, the beamforming design algorithms primarily proposed for systems without mutual coupling can be directly used for the case where the mutual coupling exists. Following the idea of using decoupling networks, [106] has proposed global optimal closed-form solutions for fully-connected and tree-connected BD-RIS-aided SISO systems, and provided the channel gain scaling law in the presence of mutual coupling. Specifically, for a SISO system with Rayleigh fading channels, [106] has derived the average power gain in the presence of mutual coupling at fully/tree-connected BD-RIS over that without mutual coupling, that is
@@
where $\mathbf{R}=\Re\{\mathbf{Z}_{II}\}^{-1}$ and $[\mathbf{Z}_{II}]_{m,m}=Z_{II}$, $\forall m\in\mathcal{M}$, and theoretically proved that $G^{\mathsf{MC}}\geq 1$ using the fact that $\mathsf{tr}(\mathbf{R})\geq\frac{M}{Z_{II}}$ and that $\mathsf{tr}(\mathbf{R}^{2})\geq\frac{M}{Z_{II}^{2}}$. This highlights that mutual coupling in fully/tree-connected BD-RIS can increase the average channel gain under Rayleigh fading channels. Not limited to SISO systems, [110] has considered more general multi-user multi-antenna systems, derived the mutual coupling aware channel model, and optimized BD-RIS under different modes to maximize the system sum-rate. Results show that not accurately capturing the mutual coupling effect at BD-RIS will cause a non-negligible sum-rate performance loss.
It is worth noting that the iterative algorithm in [104] is applicable to BD-RIS with all kinds of reciprocal architectures. Meanwhile, benefiting from the decoupling network, it is also possible to extend the analysis in [106] to other reciprocal architectures. Therefore, one interesting and feasible future research direction is to have a comprehensive performance comparison between different BD-RIS architectures in the presence of mutual coupling.
VI Benefits and Applications of BD-RIS

### v2

@@ VI-C Wideband Effect
Frequency dependence is an intrinsic phenomenon in circuits, where the response of inductance and capacitance varies with the frequency of an applied signal. Therefore, the response of RIS, i.e., its impedance/admittance/scattering matrices of the reconfigurable impedance network, is naturally frequency dependent [hu2023wideband]. In narrowband communications, since the bandwidth of signals is much less than the central frequency, it is reasonable to approximately assume the response of RIS to be frequency independent. However, when it comes to wideband communications, the frequency dependence in RIS is not negligible. To show how significant is the frequency dependence in RIS-aided wideband communication systems, there have been a few representative works on the wideband modeling, optimization, and implementation of D-RIS [li2021intelligent, zhang2021joint, wang2024wideband]. Nevertheless, these works model the wideband effect of D-RIS by characterizing the frequency dependence of the individual entries of its scattering matrix. This does not work for BD-RIS with interconnected elements, and thus coupled entries in its scattering matrix, making it difficult to accurately yet efficiently capture the frequency dependence in BD-RIS. In addition, there have been a few works on applying BD-RIS in wideband systems [soleymani2024maximizing, demir2024wideband], while they primarily assume a frequency-independent BD-RIS model, which is essentially not consistent with the behavior of BD-RIS in wideband scenarios. To establish a physically consistent BD-RIS model suitable for wideband systems, [li2025beyond] and [de2024beyond] have proposed to model the frequency dependence of the individual reconfigurable admittance component based on specific circuit designs. Below, we will briefly revisit how to model the frequency dependence in BD-RIS.
@@
Figure 32: The susceptance $\Im\{Y_{m,m^{\prime}}\}$ as a function of (a) frequency and (b) the value of susceptance at the central frequency $f_{\mathsf{c}}=2.4$ GHz with $L_{1}=2.5$ nH, $L_{2}=0.7$ nH, and $C\in[0.2,3]$ pF for a practical varactor diode [li2025beyond].
In reciprocal BD-RIS architectures, each reconfigurable admittance component can be modeled as a lumped circuit consisting of inductors $L_{1}$, $L_{2}$, and one tunable capacitor CC [^8] Here we consider a lossless model with $R=0$ in Fig. 29 to focus purely on the modeling and analysis of frequency dependence in BD-RIS.. Accordingly, the admittance of this circuit is a function of $C$ and the angular frequency $\omega=2\pi f$ for the incident signals:
@@
where $Y_{m,m}$ boils down to the admittance component $Y_{m}$ connected to ground. We observe from (98) that, the values of $Y_{m,m^{\prime}}$ at different frequencies are linked to each other by a common capacitor $C$. The relationship can be numerically illustrated in Fig. 32, based on a wideband system with central frequency $f_{\mathsf{c}}=2.4$ GHz and a practical varactor diode SMV1231-079 [li2025beyond]. Results in Fig. 32(a) show that within some practical bandwidth for wideband signals, the susceptance $\Im\{Y_{m,m^{\prime}}\}$ can be regarded as a linear function of frequency. This further implies that $\Im\{Y_{m,m^{\prime}}\}$ varies approximately linearly with the susceptance at central frequency, as illustrated in Fig. 32(b). This motivates possible simplifications to (98) to benefit BD-RIS optimization. The impact of having frequency-dependent BD-RIS matrices on the performance of a SISO orthogonal frequency division multiplexing (OFDM) system is shown in Fig. 33. Results show that, with increasing circuit complexity (more admittance components) in BD-RIS architectures, an increasing performance gap appears between taking into account the wideband modeling when designing BD-RIS or not. This is attributed to the more significant variation of BD-RIS matrices between subcarriers, so ignoring this variation will cause non-negligible performance loss.
@@
[demir2024wideband] Group-Connected
@@
The above modeling is again applicable to all reciprocal BD-RIS architectures as illustrated in Section III-D. Intuitively, the higher the circuit complexity (the larger the required number of admittance components) in BD-RIS architectures, the more significant the impact of wideband modeling at BD-RIS. In addition, [katsanos2024multi] has proposed to model and analyze the frequency dependence of non-reciprocal BD-RIS with constraint $\mathcal{T}_{\mathsf{non-diag}}$, by individually modeling the frequency-dependent scattering parameter of each reconfigurable element. Nevertheless, due to the lack of a unified circuit topology illustration as in reciprocal architectures, the wideband modeling and optimization for more general non-reciprocal BD-RIS architectures still remain unexplored. For clarity, the BD-RIS literature on wideband modeling and optimization has been summarized in Table VII.
(a) Group-connected BD-RIS with and without wideband modeling
(b) Forest-connected BD-RIS with and without wideband modeling
Figure 33: Average rate versus transmit power $P$ with BD-RIS having different reciprocal architectures ($M=48$, $\bar{M}\in\{1,3,6\}$). The OFDM channels from the transmitter to the reciever and through BD-RIS are modeled as multi-tap finite duration impulse response sequences with i.i.d. complex Gaussian random variables. The legend “WM” is short for wideband modeling; “GC” is short for group-connected; “FC” is short for forest-connected. For both “GC” and “FC” architectures, the case of $\bar{M}=1$ refers to the D-RIS [li2025beyond].
VI-D Mutual Coupling Effect
@@
Mutual coupling, which refers to the electromagnetic interaction between antenna elements in an array, is an important factor in array design. Mutual coupling is in general inversely proportional to the spacing between elements: the smaller the inter-element spacing, the stronger the mutual coupling. In most existing RIS literature, the mutual coupling is ignored by assuming sufficiently large inter-element spacing. However, this strong assumption cannot be easily achieved in practice, given that RIS usually consists of numerous densely spaced elements within a limited aperture to provide sufficient beamforming gain [di2020smart, di2022communication]. There have been existing works on the modeling and analysis of the mutual coupling between D-RIS elements [gradoni2021end, qian2021mutual, akrout2023physically], which have been recently extended to BD-RIS-aided channels [li2024beyond, semmler2024decoupling, nerini2024global]. Below, we will briefly revisit the modeling of mutual coupling aware BD-RIS-aided channels, and how the existence of mutual coupling impacts the system performance.
@@
Consider a BD-RIS-aided MIMO system consisting of an $N$-antenna transmitter, an $M$-element BD-RIS, and an $N_{\mathsf{r}}$ receiver, as also illustrated in Section IV-B. To capture explicitly the mutual coupling at BD-RIS, [universal2024, li2024beyond, gradoni2021end, akrout2023physically] have made physics-consistent assumptions and derived the following three equivalent wireless channel models [^9] Note that the scattering matrices $\mathbf{S}_{RT}$, $\mathbf{S}_{RI}$, and $\mathbf{S}_{IT}$ in (99a) are not one-to-one mapped to the channels $\mathbf{H}_{RT}$, $\mathbf{H}_{RI}$, and $\mathbf{H}_{IT}$ in the widely adopted model $\mathbf{H}(\mathbf{\Theta})=\mathbf{H}_{RT}+\mathbf{H}_{RI}\mathbf{\Theta}\mathbf{H}_{IT}$. The detailed analysis can be found in [universal2024, nossek2024physically, nerini2024physics].
@@
where $\mathbf{S}_{RT}\in\mathbb{C}^{N_{\mathsf{r}}\times N}$, $\mathbf{Z}_{RT}\in\mathbb{C}^{N_{\mathsf{r}}\times N}$, and $\mathbf{Y}_{RT}\in\mathbb{C}^{N_{\mathsf{r}}\times N}$, respectively, denote the transmission scattering, impedance, and admittance matrices from the transmitter to receiver; $\mathbf{S}_{RI}\in\mathbb{C}^{N_{\mathsf{r}}\times M}$, $\mathbf{Z}_{RI}\in\mathbb{C}^{N_{\mathsf{r}}\times M}$, and $\mathbf{Y}_{RI}\in\mathbb{C}^{N_{\mathsf{r}}\times M}$, respectively, denote the transmission scattering, impedance, and admittance matrices from RIS to the receiver; $\mathbf{S}_{IT}\in\mathbb{C}^{M\times N}$, $\mathbf{Z}_{IT}\in\mathbb{C}^{M\times N}$, and $\mathbf{Y}_{IT}\in\mathbb{C}^{M\times N}$, respectively, denote the transmission scattering, impedance, and admittance matrices from the transmitter to RIS; $\mathbf{S}_{II}\in\mathbb{C}^{M\times M}$, $\mathbf{Z}_{II}\in\mathbb{C}^{M\times M}$, and $\mathbf{Y}_{II}\in\mathbb{C}^{M\times M}$, respectively, denote the mutual coupling scattering, impedance, and admittance matrices at RIS. Based on the derivations in [universal2024] and [li2024beyond], these terms are related to one another by $\mathbf{S}_{RT}=\frac{\mathbf{Z}_{RT}}{2Z_{0}}-\frac{\mathbf{Z}_{RI}}{2Z_{0}}(\mathbf{Z}_{II}+Z_{0}\mathbf{I}_{M})^{-1}\mathbf{Z}_{IT}$, $\mathbf{S}_{RI}=\mathbf{Z}_{RI}(\mathbf{Z}_{II}+Z_{0}\mathbf{I}_{M})^{-1}$, $\mathbf{S}_{IT}=(\mathbf{Z}_{II}+Z_{0}\mathbf{I}_{M})^{-1}\mathbf{Z}_{IT}$, $\mathbf{S}_{II}=(\mathbf{Z}_{II}+Z_{0}\mathbf{I}_{M})^{-1}(\mathbf{Z}_{II}-Z_{0}\mathbf{I}_{M})$, $\mathbf{Y}_{RT}=\frac{1}{Z_{0}^{2}}(-\mathbf{Z}_{RT}+\mathbf{Z}_{RI}\mathbf{Z}_{II}^{-1}\mathbf{Z}_{IT})$, $\mathbf{Y}_{RI}=-\frac{\mathbf{Z}_{RI}\mathbf{Z}_{II}^{-1}}{Z_{0}}$, $\mathbf{Y}_{IT}=-\frac{\mathbf{Z}_{II}^{-1}\mathbf{Z}_{IT}}{Z_{0}}$, and $\mathbf{Y}_{II}=\mathbf{Z}_{II}^{-1}$.
@@
where $\mathbf{\Theta}_{\mathsf{eff}}=(\mathbf{I}_{M}-\mathbf{\Theta}\mathbf{S}_{II})^{-1}\mathbf{\Theta}$, one can also obtain a beyond-diagonal matrix $\mathbf{\Theta}_{\mathsf{eff}}$ for D-RIS. However, the BD-RIS terminology is used to describe the scattering matrix $\mathbf{\Theta}$. In this sense, within the BD-RIS, a kind of tunable coupling is artificially introduced between ports of the reconfigurable impedance network, thereby generating $\mathbf{\Theta}$ itself that is not limited to be diagonal.
Remark 10: Another interpretation of BD-RIS has been recently presented in [del2025physics], whose main idea is that an $M$-element BD-RIS-aided channel can be represented as a channel involving an effective D-RIS matrix $\tilde{\mathbf{\Theta}}\in\mathbb{C}^{\tilde{M}\times\tilde{M}}$, where $\tilde{M}$ denotes the number of reconfigurable admittance components in BD-RIS. Specifically, the channel model can be described as
@@
The impedance matrix $\mathbf{Z}_{II}$ characterizes the mismatching and mutual coupling at the RIS elements. Specifically, the diagonal entries of $\mathbf{Z}_{II}$ refer to the self impedance and the off-diagonal entries refer to the mutual coupling depending on the inter-element spacing. Most existing literature assumes perfect matching and no mutual coupling at RIS elements, which is mathematically described as $\mathbf{Z}_{II}=Z_{0}\mathbf{I}_{M}$. This assumption makes $\mathbf{S}_{II}=\mathbf{0}_{M\times M}$ and the wireless channel $\mathbf{H}_{\mathsf{S}}(\mathbf{\Theta})$ in (99) a linear function of $\mathbf{\Theta}$, as widely assumed in most existing RIS literature. However, as discussed above, it is difficult to achieve such a strong assumption in practice. That is, the off-diagonal entries of $\mathbf{Z}_{II}$ are generally nonzero, and can be modeled as functions of inter-element spacing [akrout2023physically, gradoni2021end].
@@
Figure 34: Channel gain versus the number of BD-RIS elements ($M=16$, $\bar{M}=4$ for group-connected architecture). The direct transmitter-receiver link are assumed to be fully blocked. The markers “FC”, “GC”, “SC” respectively refer to fully-connected, group-connected, and single-connected (D-RIS); “w/ MC” refers to the case accounting for mutual coupling at BD-RIS and “w/o MC” refers to the case where the off-diagonal entries of $\mathbf{Z}_{II}$ are forced to zero [li2024beyond].
@@
To evaluate how the existence of mutual coupling impacts the system performance, [li2024beyond] has primarily focused on a group/fully-connected BD-RIS-aided SISO system and proposed an iterative algorithm to design the BD-RIS impedance matrix. Simulation results in Fig. 34 demonstrate that, under purely LoS channels, the performance gap between BD-RIS and D-RIS increases with decreasing inter-element distance since the stronger mutual coupling is better exploited by BD-RIS architectures. This indicates that BD-RIS is suitable for compact deployment with small inter-element spacing. In addition, [semmler2024decoupling] has proposed decoupling networks as a solution to compensate for the mutual coupling at BD-RIS. The decoupling network can transform the system into a structure equivalent to a system without mutual coupling. As such, the beamforming design algorithms primarily proposed for systems without mutual coupling can be directly used for the case where the mutual coupling exists. Following the idea of using decoupling networks, [nerini2024global] has proposed global optimal closed-form solutions for fully-connected and tree-connected BD-RIS-aided SISO systems, and provided the channel gain scaling law in the presence of mutual coupling. Specifically, for a SISO system with Rayleigh fading channels, [nerini2024global] has derived the average power gain in the presence of mutual coupling at fully/tree-connected BD-RIS over that without mutual coupling, that is
@@
where $\mathbf{R}=\Re\{\mathbf{Z}_{II}\}^{-1}$ and $\mathbf{Z}_{II}]_{m,m}=Z_{II}$, $\forall m\in\mathcal{M}$, and theoretically proved that $G^{\mathsf{MC}}\geq 1$ using the fact that $\mathsf{tr}(\mathbf{R})\geq\frac{M}{Z_{II}}$ and that $\mathsf{tr}(\mathbf{R}^{2})\geq\frac{M}{Z_{II}^{2}}$. This highlights that mutual coupling in fully/tree-connected BD-RIS can increase the average channel gain under Rayleigh fading channels. Not limited to SISO systems, [wijekoon2024physically] has considered more general multi-user multi-antenna systems, derived the mutual coupling aware channel model, and optimized BD-RIS under different modes to maximize the system sum-rate. Results show that not accurately capturing the mutual coupling effect at BD-RIS will cause a non-negligible sum-rate performance loss. More recently, [wu2025beyondnew] has proven that band-connected architecture can achieve the same channel shaping capability as fully-connected architecture in multi-user MIMO systems with practical channel models capturing mutual coupling and other electromagnetic factors, which further generalizes the results in [nerini2024beyond, wu2025beyond, nerini2024global]. While the aforementioned works [li2024beyond, semmler2024decoupling, nerini2024global, wijekoon2024physically, wu2025beyondnew] assume idealized BD-RIS hardware to better understand the impact of mutual coupling, [zhou2025beyond] has focused on more general situations to discover optimal BD-RIS architectures that have the best performance-complexity tradeoff with mutual coupling, losses, and discrete values.
VII Applications of BD-RIS